The specific heat capacity of water is approximately 4,200 J/K/kg. Calculate the energy required to raise the temperature of 1.5 L of water from 18°C to 98°C. The density of water is 1000 kg/m3.
1
Expert's answer
2020-12-06T17:20:55-0500
"Q=mc(T_f-T_i)=\rho Vc(T_f-T_i)\\
Q=(1000)(1.5\cdot 0.001)(4200)(98-18)\\Q=504000\ J=504\ kJ"